OpenStudy (jennychan12):

Find the value(s) of c guarenteed by the Mean Value Theorem for Integrals for the function over the closed interval. f(x) = x-2x^1/2 [0,2]

OpenStudy (jennychan12):

@ChmE @agent0smith @Carl_Pham @SamuelAlden917 anyone? my teacher didn't explain this, and i'm so confused...

OpenStudy (anonymous):

\[\int\limits_{0}^{2}x-2x^{1/2}dx\]

OpenStudy (anonymous):

can you integrate this by yourself

OpenStudy (jennychan12):

yeah, thanks

OpenStudy (anonymous):

\[{x^2 \over 2}-2x^{3/2}\frac{ 2 }{ 3 }={x^2 \over 2}-{4x^{3/2} \over 3}|_{0}^{2}\]

OpenStudy (anonymous):

did you get the same thing?

OpenStudy (anonymous):

Now we have to plug in 2 for x and subtract plugging in 0 for x

OpenStudy (jennychan12):

where'd u get the 2/3 from?

OpenStudy (anonymous):

http://www.sosmath.com/calculus/integ/integ04/integ04.html 2/3 came from the 3/2 in the denominator. cuz you add one to the power all over the power. So i flipped it to change it to multiplication

OpenStudy (anonymous):

So I don't remember how to find c. I don't remember doing this type of problem. The integration is right. So now we have to work together with the link I posted to find c.

OpenStudy (jennychan12):

yeah i got the F(x) same thing..

OpenStudy (anonymous):

Ok i got it

OpenStudy (anonymous):

http://www.dummies.com/how-to/content/using-the-mean-value-theorem-for-integrals.html Scroll down to the bottom

OpenStudy (jennychan12):

ok i think i got it.

OpenStudy (anonymous):

\[f(c)=\frac{ 1 }{ 2-0 }[({(2)^2 \over 2}-{4(2)^{3/2} \over 3})-({(0)^2 \over 2}-{4(0)^{3/2} \over 3})]\]

OpenStudy (anonymous):

Post your answer I'll check it with mine

OpenStudy (jennychan12):

yeah i got the same thing on paper.

OpenStudy (anonymous):

my final answer is\[1-\frac{ 4\sqrt2 }{ 3 }\]

OpenStudy (jennychan12):

i got \[2-\frac{ 8\sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

don't forget about distributing the 1/2

OpenStudy (jennychan12):

whoops i forgot to multiply by 1/2

OpenStudy (anonymous):

NIce job. I learned something tonight too

OpenStudy (jennychan12):

wait... there are two answers..... 0.4380, 1.7908

OpenStudy (anonymous):

Is that what the book says?

OpenStudy (jennychan12):

yeah

OpenStudy (anonymous):

I don't see how

OpenStudy (anonymous):

either of those answers don't match ours

OpenStudy (jennychan12):

this is what the answer thingy says

OpenStudy (jennychan12):

i think i can try to figure it out.

OpenStudy (anonymous):

I see it

OpenStudy (jennychan12):

sorry i'm a little lost...

OpenStudy (anonymous):

We got our f(c) right. \[f(x) = x-2x^{1/2}\]\[f(c)=1-\frac{ 4\sqrt2 }{ 3 }\]\[c-2c^{1/2}=1-\frac{ 4\sqrt2 }{ 3 }\]Solve for c

OpenStudy (anonymous):

They entered c in for x in the original formula and set it equal to f(c)

OpenStudy (jennychan12):

ohh i see it.

OpenStudy (anonymous):

The next step looks like they completed the square

OpenStudy (anonymous):

Then it gets really messy

OpenStudy (anonymous):

At least you know the basics of how to find c now

OpenStudy (jennychan12):

ok thanks.